OpenStudy (anonymous):

combine as indicated by the sign : (x+2)/(x+4) - (x-1)/(x+6)

OpenStudy (anonymous):

Any problems? Hint: Common denominator

OpenStudy (anonymous):

(x+2)/(x+4) - (x-1)/(x+6) 1- conman denominator (x+6)(x+4) [(x+2)(x+6)-(x-1)(x+4)]/[x+6)(x+4)

OpenStudy (anonymous):

what? the answer has a few given sections such as __x______/___6____4_

OpenStudy (anonymous):

@Lyralei I have no clue how to do this at all, any help will do please.

OpenStudy (anonymous):

\[\frac{x+2}{x+4}-\frac{x-1}{x+6}\] Whenever you want to mix a fraction, you should multiply by each other's denominator, as such: \[\frac{(x+2)(x+6)}{(x+4)(x+6)}-\frac{(x-1)(x+4)}{(x+6)(x+4)}\]

OpenStudy (anonymous):

\[\frac{x^2+8+12}{x^2+10+24}-\frac{x^2+3x-4}{x^2+10x+24}\] Notice that you could combine the fractions together now because the denominators are the same.

OpenStudy (anonymous):

\[=\frac{(x^2+8x+12)-(x^2+3x-4)}{x^2+10+24}\] \[=\frac{5x+16}{x^2+10+24}\]

OpenStudy (anonymous):

@Agreenleaf

OpenStudy (anonymous):

that is the same answer i got, but it doesnt fit with the parts of the answer that were given to me.

OpenStudy (anonymous):

ok that picture didnt work. but i was given ___x______ / ___6____4_ where the lines are blank spots i need to fill out @Lyralei

OpenStudy (anonymous):

I know what it is, you'll have to fill out the blanks, we have the numerator 5x+16, but the 6 and 4 in the denominator is actually in factorised form.

OpenStudy (anonymous):

Like this \[\frac{5 x+16}{(x+4) (x+6)}\]

OpenStudy (anonymous):

omg that helps haha, thanks!

OpenStudy (anonymous):

No problem!